# Ramification filtration via deformations, II

### v3

@@ 4.7. Lie elements ¯$\bar{\mathcal{F}}[\iota]$ and ¯$\bar{\mathcal{F}}[\iota]_{0}$
The following proposition relies elements ¯$\bar{\mathcal{F}}[\iota]$ and the elements ¯$\bar{\mathcal{F}}_{\gamma,-N}$ from Section 4.3.
@@
If $\iota$ is given in notation (4.1) then
@@
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\eta(1,\dots,s)a_{1}p^{n_{1}}[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{s}n_{s}}]=\bar{\mathcal{F}}^{0}_{\gamma(iota),-(M(\iota)+n)}\,.$
@@
Corollary 4.8.
All ¯$\bar{\mathcal{F}}[\iota p^{m}]$ and ¯$\bar{\mathcal{F}}[\iota p^{m}]_{0}$ belong to ¯$\bar{\mathcal{L}}_{W_{M}(k)}$.
This finishes the proof of Theorem 4.1.
4.8. Description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ in [3]
Now we use Theorem 4.1 to recover the description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ obtained in [3].
@@ 4.8. Description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ in [3]
Let $\widetilde{N}=N^{*}-1$. Then the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}_{W_{M}(k)}^{[v_{0}]}$ contains all ¯$\bar{\mathcal{F}}^{0}_{\alpha,-\widetilde{N}}$ with $\gamma\geqslant v_{0}$.
@@
By Theorem 4.1, ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$ contains all ¯$\bar{\mathcal{F}}[\iota p^{r_{\iota}}]$, where $\iota\in\mathfrak{A}_{1}^{+}(p)$ and $r_{\iota}\geqslant 0$ is such that $\iota p^{r_{\iota}+1}\notin\mathfrak{A}^{0}$.
@@
$\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})=p^{\widetilde{M}(\iota)+M-1}(q\gamma(\iota)-b^{*})\in\mathfrak{A}_{1}^{+}(p)\,,$
where $\gamma(\iota)=p^{A}\alpha$, $\alpha\in A[p-1,m]$ and $\widetilde{M}(\iota)=m-A$. Note that $\iota\in\mathfrak{A}^{+}_{1}(p)$ means also that $\iota\leqslant p^{M-1}(p-1)b^{*}$ and we have the following two cases (cf. Prop. 4.10 below): either $v_{p}(\gamma(\iota)+\widetilde{M}\iota)=0$ or $\widetilde{M}(\iota)+M-1=0$ (what is equivalent to $A=M-1$ amd $m=0$).
Then by Prop. 4.7, ¯$\bar{\mathcal{F}}[\iota p^{r_{\iota}}]$ can be replaced by ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-\widetilde{N}}$ because $\widetilde{N}=N^{*}-1=\widetilde{M}(\iota)+r_{\iota}$ for some $r_{\iota}\geqslant r(\iota)$, where as earlier
$r(\iota)=\max\{r\geqslant 0\ |\ \iota p^{r(\iota)}\in\mathfrak{A}^{0}\}\,.$
@@
Suppose $\gamma\geqslant v_{0}$ is such that for any $\iota\in\mathfrak{A}^{+}_{1}(p)$, $\gamma\neq\gamma(\iota)$ and $v_{p}(\gamma)\geqslant-\widetilde{N}$. Then ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$.
@@
As it was mentioned above, we have the two following cases:
1) $p^{M-1-v_{p}(\gamma)}(q\gamma-b^{*})>p^{M-1}(p-1)b^{*}$;
2) $v_{p}(\gamma)\geqslant 0$ and $q\gamma-b^{*}>p^{M-1}(p-1)b^{*}$.
@@
where $M>n_{1}\geqslant n_{2}\geqslant\ldots\geqslant n_{r}\geqslant-\widetilde{N}$ and $p^{n_{1}}a_{1}+\dots+p^{n_{r}}a_{r}=\gamma$.
We can assume that all $a_{i}<(p-1)v_{0}$, because otherwise ${\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\mathcal{L}(p)_{W_{M}(k)}$ and ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}=0$.
Consider the first case.
From definitions it follows that ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}$ is a linear combination of the terms of the form
@@
If $a_{1}^{\prime}+\dots+a_{r}^{\prime}\geqslant(p-1)b^{*}/q$, i.e. it is $\geqslant(p-1)v_{0}$, then the whole term belongs to ¯$\bar{\mathcal{L}}(p)_{W_{M}(k)}=0$.
Otherwise, the inequality
@@
As a result we either prove our proposition or will come to the second case.
Here we proceed similarly using the above calculations with $\widetilde{N}=0$:
$\gamma=\gamma^{\prime}+(a_{1}^{\prime}+\ldots+a^{\prime}_{r})$, $r\geqslant 0$ and $v_{p}(\gamma^{\prime})\geqslant 1$. Then
$q\gamma-b^{*}=q\gamma^{\prime}-b^{*}+q(a_{1}^{\prime}+\ldots+a^{\prime}_{r})>p^{M-1}(p-1)b^{*}\,.$
Again, if $a_{1}^{\prime}+\ldots+a^{\prime}_{r}\geqslant(p-1)v_{0}$ we are in ¯$\bar{\mathcal{L}}(p)_{W_{M}(k)}$. Otherwise, $q\gamma^{\prime}-b^{*}>0$ implies $\gamma^{\prime}\geqslant v_{0}$ and we have alternative: $\gamma^{\prime}=\gamma(\iota^{\prime})$ with $\iota^{\prime}\in\mathfrak{A}_{1}^{+}(p)$ or repeat the procedure. ∎

### v5

@@ 4.7. Lie elements ¯$\bar{\mathcal{F}}[\iota]$ and ¯$\bar{\mathcal{F}}[\iota]_{0}$
The following proposition relates the elements ¯$\bar{\mathcal{F}}[\iota]$ and the elements ¯$\bar{\mathcal{F}}_{\gamma,-N}$ from Section 4.3.
@@
If $\iota$ is given in notation (4.1) and $n\geqslant 0$ then
@@
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\eta(1,\dots,s)a_{1}p^{n_{1}}[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{s}n_{s}}]=\bar{\mathcal{F}}^{0}_{\gamma(\iota),-(M(\iota)+n)}\,.$
@@
According to remark from Section 4.5 this finishes the proof of Theorem 4.1 (we should use $n=m_{\iota}$ in the last Proposition).
4.8. Description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ from [3]
We are going to use Theorem 4.1 to recover the description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ obtained in [3] in the context of Lie algebra $\bar{\mathcal{L}}$.
@@ 4.8. Description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ from [3]
Let $\widetilde{N}=N^{*}-1$. Then the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}_{W_{M}(k)}^{[v_{0}]}$ contains all ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}$ with $\gamma\geqslant v_{0}$.
@@
By Theorem 4.1, ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W(k)}$ contains all ¯$\bar{\mathcal{F}}[\iota p^{m_{\iota}}]$, where $\iota\in\mathfrak{A}_{1}^{+}(p)$ and $m_{\iota}\geqslant 0$ is chosen such that $\iota p^{m_{\iota}}\notin\mathfrak{A}^{0}$.
@@
$\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})=p^{M(\iota)+M-1}(q\gamma(\iota)-b^{*})\in\mathfrak{A}_{1}^{+}(p)\,,$
where $\gamma(\iota)=p^{A}\alpha$, $\alpha\in A[p-1,m]$ and $M(\iota)=m-A$. Note that $\iota\in\mathfrak{A}^{+}_{1}(p)$ means mainly that $\iota\leqslant p^{M-1}(p-1)b^{*}$. In this case by Prop. 4.5, ¯$\bar{\mathcal{F}}[\iota p^{m_{\iota}}]$ can be replaced by ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-\widetilde{N}}$ because for some $m_{\iota}\geqslant r(\iota)$, $\widetilde{N}=N^{*}-1={M}(\iota)+m_{\iota}$. Note also that $\gamma(\iota)\in A[p-1,M(\iota)]$.
@@
Suppose $\gamma\geqslant v_{0}$ is such that for any $\iota\in\mathfrak{A}^{+}_{1}(p)$, $\gamma\neq\gamma(\iota)$ and $\gamma\in A[p-1,\widetilde{N}]$. Then ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\bar{\mathcal{L}}^{[v_{0}]}_{W(k)}$.
@@
We can assume that $p^{M-1+\widetilde{N}}(q\gamma-b^{*})>p^{M-1}(p-1)b^{*}\,.$
@@
where $M>n_{1}\geqslant n_{2}\geqslant\ldots\geqslant n_{r}\geqslant-\widetilde{N}$ and $p^{n_{1}}a_{1}+\dots+p^{n_{r}}a_{r}=\gamma$. We can assume that all $a_{i}<(p-1)v_{0}$, because otherwise ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\bar{\mathcal{L}}(p)_{W(k)}$ and ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}=0$.
Note that ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}$ can be presented as a linear combination of the terms
@@
If $a_{1}^{\prime}+\dots+a_{r}^{\prime}\geqslant(p-1)b^{*}/q$, i.e. it is $\geqslant(p-1)v_{0}$, then the whole term belongs to ¯$\bar{\mathcal{L}}(p)_{W(k)}=0$. Otherwise, the inequality
@@
As a result we either prove our proposition or will come to the case where we have ¯$\bar{\mathcal{F}}_{\gamma,0}$ with $q\gamma-b^{*}>p^{M-1}(p-1)b^{*}$. Proceed as earlier:
Let $\gamma=\gamma^{\prime}+(a_{1}^{\prime}+\ldots+a^{\prime}_{r})$, $r\geqslant 0$ and $v_{p}(\gamma^{\prime})\geqslant 1$. Then
$q\gamma-b^{*}=q\gamma^{\prime}-b^{*}+q(a_{1}^{\prime}+\ldots+a^{\prime}_{r})>(p-1)b^{*}\,.$
Again, if $a_{1}^{\prime}+\ldots+a^{\prime}_{r}\geqslant(p-1)v_{0}$ we are in ¯$\bar{\mathcal{L}}(p)_{W(k)}$. Otherwise, $q\gamma^{\prime}-b^{*}>0$ implies $\gamma^{\prime}\geqslant v_{0}$ and we have the following alternative: $\gamma^{\prime}=\gamma(\iota^{\prime})$ with $\iota^{\prime}\in\mathfrak{A}_{1}^{+}(p)$ or should repeat the procedure with ¯$\bar{\mathcal{F}}_{\gamma^{\prime},1}\in p\bar{\mathcal{L}}_{W(k)}$. It remains to note that ¯$\bar{\mathcal{F}}_{\gamma,M}=0$. ∎